# Splitting criteria for ordinal decision trees: an experimental study

### v1

@@ Abstract
I Introduction
@@
OC problems can be found in a wide variety of fields, including aforementioned medical research [5, 6, 7, 8, 9], automation science and engineering [10], wave height prediction [11], credit rating [12], face recognition [13, 14, 15], social events forecasting [16], age estimation [17], image classification [18], wind speed prediction [19, 20], manufacturing [21], text classification [22], and transportation [23], among others. These studies exemplify the use of specifically designed OC models, which leverage the ordinal information to enhance performance compared to their nominal counterparts.
The OC paradigm has considerably evolved in recent years, thanks to substantial advances in supervised learning methodologies [24]. This includes contributions ranging from deep regression models [25] or discriminant learning techniques [26], to graph-based approaches [27] and soft-labelling methods [28], among others. Tree-based models for OC have also been a hot topic in the last decade [29, 30, 31, 32]. One of the primary advantages of tree-based approaches is their ability to produce interpretable models that facilitate the understanding of complex and non monotonic dependencies between explanatory attributes and the dependent variable, also referred as output variable. For instance, the simplest tree-based models, the Decision Trees (DTs), can generate easy-to-interpret rules without compromising performance. Another strength of DTs is their versatility, as they can handle continuous, discrete and categorical inputs while being invariant under strictly monotone transformations of individual inputs. In addition, an internal feature selection is performed as an integral part of the training process.
@@
Specifically, the Gini-index impurity measure computes the expected loss on all classes due to misclassification errors, whereas the Shannon-entropy represents the uncertainty of a random variable outcome. However, both Gini-index and Shannon-entropy, to be minimised, do not take into account the ordinal properties of the dependent variable. For instance, lets suppose the following two nodes representing the distribution of $20$ observations into $4$ ordered classes: A) $(10,0,0,10)$, and B) $(10,10,0,0)$. Considering these two distributions, we can observe that both Gini-index and Shannon-entropy result in the same value ($0.5$ and $1.0$, respectively). However, from an ordinal perspective, option B is preferable over option A, as it better separates the classes following an ordinal scale. This example is graphically illustrated in Figure 2, where the values for ordinal impurity measures are also presented (detailed along with their associated splitting criterion in Section III). This observation extends to the splitting criteria, to be maximised, defined by these impurity measures. Figure 3 illustrates two different splits of a given node, for which the nominal splitting criteria achieve the same value regardless the ordinality considered in the sub-nodes.
@@
The rest of the paper is organised as follows: Section II reviews commonly used impurity measures (Section II-A) and main OC tree-based approaches (Section II-B); Section III presents the nominal and ordinal impurity measures and their corresponding splitting criteria; Section IV describes the experimental study conducted; the results achieved are included and statistically analysed in Section V; and finally, Section VI closes this work with some final remarks and future research directions.
II Literature review
@@ II-A Impurity measures
As observed, most of the aforementioned ordinal impurity measures depend on the assignment of numerical scores to the classes, thereby using a higher scale level. Specifically, the $H_{w}$ and RI explicitly incorporate these assigned scores into their definitions. In contrast, OGini-index is score-free as its definition is based on the cumulative frequency of each class. While the assignment of scores may be justified in certain scenarios, particularly when ordinal labels are built from continuous variables by discretisation, it is rather artificial and arbitrary in the context of genuine ordinal data. For instance, when the response variable represents ordered levels of disease severity, assigning numerical scores might not align with the natural characteristics of the data. In this work, we use the three previous ordinal impurity measures to build different splitting criteria, which are thoroughly defined and mathematically formalised in Section III.
II-B Ordinal classification trees
@@
However, some approaches do not properly fit into the categories defined by either taxonomy, such as the tree-based approaches, in which a number of methods have been proposed in recent years. Beyond the approaches detailed in Section II-A, additional methods have been introduced in the literature, such as the one proposed in [4], which is based on the concept of consistency, a global property that characterises the relationship between different decision regions within the input space. This property was integrated into the design of the learning algorithm, where it guides the node expansion process to ensure alignment with the ordinal structure of the data.
@@ III-A Gini
III-B Ordinal Gini (OGini)
@@ III-C Information Gain (IG)
III-D Weighted Information Gain (WIG)
@@ III-E Ranking Impurity (RI)
IV Experimental setup
This section describes the experimental framework used to evaluate the differences in terms of performance for the various splitting criteria presented in Section III. These splitting criteria include Gini, OGini, IG, WIG, and RI, which are utilised to build a DT classifier. Note that all other hyperparameters of the DT classifier remain fixed, with each DT classifier being named after the splitting criterion used during its training phase.
To conduct a comprehensive study, we evaluate each splitting criterion using a set of $45$ ordinal datasets collected from different OC works. Specifically, $41$ of them were obtained from [45], which is the the most extensive experimental study conducted to date in the OC field. The remaining $4$ datasets are well-established and frequently utilised [30]. A detailed overview of these datasets is provided in Table I.
@@ IV Experimental setup
TABLE I: Characteristics of the datasets used in this work. #Train and #Test stand for the number of training and testing patterns, respectively, $Q$ indicates the number of classes, and $K$ is the number of input features.
Dataset #Train #Test $Q$ $K$
Original OC problems ERA $750$ $250$ $9$ $4$
ESL $366$ $122$ $9$ $4$
LEV $750$ $250$ $5$ $4$
SWD $750$ $250$ $4$ $10$
automobile $153$ $52$ $6$ $71$
balance-scale $468$ $157$ $3$ $4$
bondrate $42$ $15$ $4$ $37$
car $1296$ $432$ $4$ $21$
contact-lenses $18$ $6$ $3$ $6$
eucalyptus $552$ $184$ $5$ $91$
mammoexp $276$ $136$ $3$ $5$
newthyroid $161$ $54$ $3$ $5$
nhanes $3499$ $1724$ $5$ $30$
pasture $27$ $9$ $3$ $25$
squash-stored $39$ $13$ $3$ $51$
squash-unstored $39$ $13$ $3$ $52$
support $489$ $241$ $5$ $19$
tae $113$ $38$ $3$ $54$
toy $225$ $75$ $5$ $2$
vlbw $115$ $57$ $9$ $19$
winequality-red $1199$ $400$ $6$ $11$
Discretised regression problems abalone $1000$ $3177$ $5$ $10$
abalone10 $1000$ $3177$ $10$ $10$
bank1-10 $50$ $8142$ $10$ $8$
bank1-5 $50$ $8142$ $5$ $8$
bank2-10 $75$ $8117$ $10$ $32$
bank2-5 $75$ $8117$ $5$ $32$
calhousing-10 $150$ $20490$ $10$ $8$
calhousing-5 $150$ $20490$ $5$ $8$
census1-10 $175$ $22609$ $10$ $8$
census1-5 $175$ $22609$ $5$ $8$
census2-10 $200$ $22584$ $10$ $16$
census2-5 $200$ $22584$ $5$ $16$
computer1-10 $100$ $8092$ $10$ $12$
computer1-5 $100$ $8092$ $5$ $12$
computer2-10 $125$ $8067$ $10$ $21$
computer2-5 $125$ $8067$ $5$ $21$
housing $300$ $206$ $5$ $13$
housing10 $300$ $206$ $10$ $13$
machine $150$ $59$ $5$ $6$
machine10 $150$ $59$ $10$ $6$
pyrim $50$ $24$ $5$ $26$
@@
V Results
The results obtained for the $20$ executions of each DT classifier using different splitting criteria are presented in Table II, expressed as $\text{Mean}_{\text{STD}}$, where STD stands for the standard deviation. It can be observed that the ordinal splitting criteria, i.e. OGini, WIG, and RI, obtain better results for MAE and QWK than the nominal ones (Gini and IG). In terms of RPS, OGini and RI again outperform IG and Gini. However, in this case WIG fails to outperform the nominal metrics. It is important to remark that OGini obtains the best mean results for the three metrics, while the second best results are obtained by the WIG in MAE and QWK, and by the RI in terms of RPS.
TABLE II: Mean and standard deviation (${}_{\text{STD}}$) for each method in terms of MAE, QWK and RPS. All datasets and splits are considered.
@@ V Results
V-A Statistical analysis for datasets with $Q\geq 6$
In terms of MAE, the null hypothesis of normality of the K-S test is accepted for each of the $85$ samples of size $20$ ($17$ databases × $\times\ 5$ splitting criteria). Subsequently, the ANOVA II test is performed to test whether the methodology and dataset (factors) have any significant impact on the mean values obtained in each metric. The results indicate the existence of significant differences between the mean results obtained by the different methodologies (p-value $<0.001$). In view of these results, the Tukey post-hoc test is conducted. The results for this test are presented in Table III, where the methods (splitting criteria) are grouped based on the statistical differences in their performance. Methods within the same group exhibit no significant differences, while those in different groups show statistically significant differences. As can be observed in the left part of Table III ($Q\geq 6$), group $1$ contains the best methodologies, OGini and RI, with no significant differences between them (p-value $=0.063$), and the group $3$ contains the worst methodologies, Gini and IG. This table also indicates significant differences between the pairs OGini-Gini, OGini-IG, WIG-IG, and RI-IG, always in favour of the ordinal methods. However, although RI presents a lower MAE, the difference against Gini is not significant (group $2$). Additionally, OGini stands as the best performing splitting criterion, obtaining significant differences against all the other criteria except for the RI.
TABLE III: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of MAE.
@@ V-A Statistical analysis for datasets with $Q\geq 6$
For the QWK metric, the K-S test accepts the null hypothesis of normality, so the same procedure utilised for the MAE is conducted. The ANOVA II shows an overall statistically significant difference between the mean results across the splitting criteria (p-value $<0.001$). Therefore, the post-hoc Tukey test is conducted (Table IV). OGini and RI outperform both IG and Gini, and WIG surpasses IG. However, WIG shows no significant differences compared to Gini (p-value $=0.089$). Similarly as for the MAE metric, OGini stands as the best approach, having no significant differences against RI and WIG (p-value $=0.996$).
TABLE IV: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of QWK.
@@
Finally, in terms of RPS, once the hypothesis of normality is accepted, the ANOVA II test demonstrate a significant difference between the mean results of the five splitting criteria (p-value $<0.001$). The Tukey test (see Table V) indicates that there are no significant differences in mean RPS between RI, OGini, Gini and IG (p-value $=0.082$). On the contrary, WIG is observed to be significantly worse than RI, Ogini and Gini. Additionally, we can observe that RI obtains the best results.
TABLE V: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of RPS.
@@
As a general comment, by comparing left and right parts of Tables III, IV and V, it is clear that differences are more evident in performance when the number of classes increases. This comes from the fact that the possible costs of misclassification increases, and ordinal splitting criteria can better exploit the ordinality of the target label and its relationship with the input data.
V-B Statistical analysis for datasets with $Q<6$
Focusing on the MAE metric, the hypothesis of normality is accepted, then, the ANOVA II is applied, showing that there are significant differences between the mean values of the five splitting criteria (p-value$=0.049$). Thus, the post-hoc Tukey test is conducted. The results are shown in the right part of Table III. It can be observed that OGini achieves the best mean results. However, there are no significant differences between OGini, RI, and WIG with respect to Gini and IG.
In terms of QWK, since the hypothesis of normality is accepted, the ANOVA II test is applied indicating that no significant differences are found with respect to the mean results obtained by the splitting criteria, achieving a p-value$=0.446$. In this case, the post-hoc Tukey test is not performed (in Table IV all the splitting criteria are in the same group).
Finally, for the RPS metric the hypothesis of normality is accepted again. Then, the ANOVA II test is conducted, showing significant differences in the mean values between the different splitting criteria. Hence, the post-hoc Tukey test is performed. The results (see Table V) indicate that OGini is the best splitting criteria in mean results, but no significant differences are found with respect to Gini and IG.
V-C Analysis of the confusion matrices
@@
TABLE VI: Confusion matrices obtained in the test set by Gini and OGini in the census-2-5 dataset together with the QWK, MAE and RPS values.

### v2

@@ Abstract
journal: Pattern Recognition
[…]
[uco]organization=Department of Computer Science and Numerical Analysis, Universidad de Córdoba, addressline=Campus de Rabanales, Ctra. N-IVa, Km. 396, city=Córdoba, postcode=14071, country=Spain
[…]
[uloyola]organization=Department of Quantitative Methods, Universidad Loyola Andalucía, addressline=C. Escritor Castilla Aguayo, 4, city=Córdoba, postcode=14004, country=Spain
1 Introduction
@@
OC problems can be found in a wide variety of fields, including aforementioned medical research [5, 6, 7, 8, 9], automation science and engineering [10], wave height prediction [11], credit rating [12], face recognition [13, 14, 15], age estimation [16], image classification [17], wind speed prediction [18, 19], manufacturing [20], text classification [21], and transportation [22], among others. These studies exemplify the use of specifically designed OC models, which leverage the ordinal information to enhance performance compared to their nominal counterparts.
The OC paradigm has considerably evolved in recent years, thanks to substantial advances in supervised learning methodologies [23]. This includes contributions ranging from neural network models, like the one proposed in [24], which also addresses the imbalance problem typical of ordinal tasks by minimizing a Bayesian cost, to deep ordinal models [25], soft-labelling methods [26, 27] and ordinal representation learning techniques [28]. Tree-based models for OC have also been a hot topic in the last decade [29, 30, 31, 32]. One of the primary advantages of tree-based approaches is their ability to produce interpretable models that facilitate the understanding of complex and non monotonic dependencies between explanatory attributes and the dependent variable, also referred as output variable. For instance, the simplest tree-based models, the Decision Trees (DTs), can generate easy-to-interpret rules without compromising performance. Another strength of DTs is their versatility, as they can handle continuous, discrete and categorical inputs while being invariant under strictly monotone transformations of individual inputs. In addition, an internal feature selection is performed as an integral part of the training process.
@@
Specifically, the Gini-index impurity measure computes the expected loss on all classes due to misclassification errors, whereas the Shannon-entropy represents the uncertainty of a random variable outcome. However, both Gini-index and Shannon-entropy, to be minimised, do not take into account the ordinal properties of the dependent variable. For instance, lets suppose the following two nodes representing the distribution of $20$ observations into $4$ ordered classes: A) $(10,0,0,10)$, and B) $(10,10,0,0)$. Considering these two distributions, we can observe that both Gini-index and Shannon-entropy result in the same value ($0.5$ and $1.0$, respectively). However, from an ordinal perspective, option B is preferable over option A, as it better separates the classes following an ordinal scale. This example is graphically illustrated in Figure 2, where the values for ordinal impurity measures are also presented (detailed along with their associated splitting criterion in Section 3). This observation extends to the splitting criteria, to be maximised, defined by these impurity measures. Figure 3 illustrates two different splits of a given node, for which the nominal splitting criteria achieve the same value regardless the ordinality considered in the sub-nodes.
@@
The rest of the paper is organised as follows: Section 2 reviews commonly used impurity measures (Section 2.1) and main OC tree-based approaches (Section 2.2); Section 3 presents the nominal and ordinal impurity measures and their corresponding splitting criteria; Section 4 describes the experimental study conducted; the results achieved are included and statistically analysed in Section 5; and finally, Section 6 closes this work with some final remarks and future research directions.
2 Literature review
@@ 2.1 Impurity measures
As observed, most of the aforementioned ordinal impurity measures depend on the assignment of numerical scores to the classes, thereby using a higher scale level. Specifically, the $H_{w}$ and RI explicitly incorporate these assigned scores into their definitions. In contrast, OGini-index is score-free as its definition is based on the cumulative frequency of each class. While the assignment of scores may be justified in certain scenarios, particularly when ordinal labels are built from continuous variables by discretisation, it is rather artificial and arbitrary in the context of genuine ordinal data. For instance, when the response variable represents ordered levels of disease severity, assigning numerical scores might not align with the natural characteristics of the data. In this work, we use the three previous ordinal impurity measures to build different splitting criteria, which are thoroughly defined and mathematically formalised in Section 3.
2.2 Ordinal classification trees
@@
However, some approaches do not properly fit into the categories defined by either taxonomy, such as the tree-based approaches, in which a number of methods have been proposed in recent years. Beyond the approaches detailed in Section 2.1, additional methods have been introduced in the literature, such as the one proposed in [4], which is based on the concept of consistency, a global property that characterises the relationship between different decision regions within the input space. This property was integrated into the design of the learning algorithm, where it guides the node expansion process to ensure alignment with the ordinal structure of the data.
@@ 3.1 Gini
3.2 Ordinal Gini (OGini)
@@ 3.3 Information Gain (IG)
3.4 Weighted Information Gain (WIG)
@@ 3.5 Ranking Impurity (RI)
4 Experimental setup
This section describes the experimental framework used to evaluate the differences in terms of performance for the various splitting criteria presented in Section 3. These splitting criteria include Gini, OGini, IG, WIG, and RI, which are utilised to build a DT classifier. Note that all other hyperparameters of the DT classifier remain fixed, with each DT classifier being named after the splitting criterion used during its training phase.
To conduct a comprehensive study, we evaluate each splitting criterion using a set of $45$ ordinal datasets collected from different OC works. Specifically, $41$ of them were obtained from [45], which is the the most extensive experimental study conducted to date in the OC field. The remaining $4$ datasets are well-established and frequently utilised [30]. A detailed overview of these datasets is provided in Table 1.
@@ 4 Experimental setup
Table 1: Characteristics of the datasets used in this work. #Train and #Test stand for the number of training and testing patterns, respectively, $Q$ indicates the number of classes, and $K$ is the number of input features.
Original OC problems Discretised regression problems
Dataset #Train #Test $Q$ $K$ Dataset #Train #Test $Q$ $K$
ERA $750$ $250$ $9$ $4$ abalone $1000$ $3177$ $5$ $10$
ESL $366$ $122$ $9$ $4$ abalone10 $1000$ $3177$ $10$ $10$
LEV $750$ $250$ $5$ $4$ bank1-10 $50$ $8142$ $10$ $8$
SWD $750$ $250$ $4$ $10$ bank1-5 $50$ $8142$ $5$ $8$
automobile $153$ $52$ $6$ $71$ bank2-10 $75$ $8117$ $10$ $32$
balance-scale $468$ $157$ $3$ $4$ bank2-5 $75$ $8117$ $5$ $32$
bondrate $42$ $15$ $4$ $37$ calhousing-10 $150$ $20490$ $10$ $8$
car $1296$ $432$ $4$ $21$ calhousing-5 $150$ $20490$ $5$ $8$
contact-lenses $18$ $6$ $3$ $6$ census1-10 $175$ $22609$ $10$ $8$
eucalyptus $552$ $184$ $5$ $91$ census1-5 $175$ $22609$ $5$ $8$
mammoexp $276$ $136$ $3$ $5$ census2-10 $200$ $22584$ $10$ $16$
newthyroid $161$ $54$ $3$ $5$ census2-5 $200$ $22584$ $5$ $16$
nhanes $3499$ $1724$ $5$ $30$ computer1-10 $100$ $8092$ $10$ $12$
pasture $27$ $9$ $3$ $25$ computer1-5 $100$ $8092$ $5$ $12$
squash-stored $39$ $13$ $3$ $51$ computer2-10 $125$ $8067$ $10$ $21$
squash-unstored $39$ $13$ $3$ $52$ computer2-5 $125$ $8067$ $5$ $21$
support $489$ $241$ $5$ $19$ housing $300$ $206$ $5$ $13$
tae $113$ $38$ $3$ $54$ housing10 $300$ $206$ $10$ $13$
toy $225$ $75$ $5$ $2$ machine $150$ $59$ $5$ $6$
vlbw $115$ $57$ $9$ $19$ machine10 $150$ $59$ $10$ $6$
winequality-red $1199$ $400$ $6$ $11$ pyrim $50$ $24$ $5$ $26$
@@
5 Results
The results obtained for the $20$ executions of each DT classifier using different splitting criteria are presented in Table 2, expressed as $\text{Mean}_{\text{STD}}$, where STD stands for the standard deviation. It can be observed that the ordinal splitting criteria, i.e. OGini, WIG, and RI, obtain better results for MAE and QWK than the nominal ones (Gini and IG). In terms of RPS, OGini and RI again outperform IG and Gini. However, in this case WIG fails to outperform the nominal metrics. It is important to remark that OGini obtains the best mean results for the three metrics, while the second best results are obtained by the WIG in MAE and QWK, and by the RI in terms of RPS.
Table 2: Mean and standard deviation (${}_{\text{STD}}$) for each method in terms of MAE, QWK and RPS. All datasets and splits are considered.
@@ 5 Results
5.1 Statistical analysis for datasets with $Q\geq 6$
In terms of MAE, the null hypothesis of normality of the K-S test is accepted for each of the $85$ samples of size $20$ ($17$ databases × $\times\ 5$ splitting criteria). Subsequently, the ANOVA II test is performed to test whether the methodology and dataset (factors) have any significant impact on the mean values obtained in each metric. The results indicate the existence of significant differences between the mean results obtained by the different methodologies (p-value $<0.001$). In view of these results, the Tukey post-hoc test is conducted. The results for this test are presented in Table 3, where the methods (splitting criteria) are grouped based on the statistical differences in their performance. Methods within the same group exhibit no significant differences, while those in different groups show statistically significant differences. As can be observed in the left part of Table 3 ($Q\geq 6$), group $1$ contains the best methodologies, OGini and RI, with no significant differences between them (p-value $=0.063$), and the group $3$ contains the worst methodologies, Gini and IG. This table also indicates significant differences between the pairs OGini-Gini, OGini-IG, WIG-IG, and RI-IG, always in favour of the ordinal methods. However, although RI presents a lower MAE, the difference against Gini is not significant (group $2$). Additionally, OGini stands as the best performing splitting criterion, obtaining significant differences against all the other criteria except for the RI.
Table 3: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of MAE.
@@ 5.1 Statistical analysis for datasets with $Q\geq 6$
For the QWK metric, the K-S test accepts the null hypothesis of normality, so the same procedure utilised for the MAE is conducted. The ANOVA II shows an overall statistically significant difference between the mean results across the splitting criteria (p-value $<0.001$). Therefore, the post-hoc Tukey test is conducted (Table 4). OGini and RI outperform both IG and Gini, and WIG surpasses IG. However, WIG shows no significant differences compared to Gini (p-value $=0.089$). Similarly as for the MAE metric, OGini stands as the best approach, having no significant differences against RI and WIG (p-value $=0.996$).
Table 4: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of QWK.
@@
Finally, in terms of RPS, once the hypothesis of normality is accepted, the ANOVA II test demonstrate a significant difference between the mean results of the five splitting criteria (p-value $<0.001$). The Tukey test (see Table 5) indicates that there are no significant differences in mean RPS between RI, OGini, Gini and IG (p-value $=0.082$). On the contrary, WIG is observed to be significantly worse than RI, Ogini and Gini. Additionally, we can observe that RI obtains the best results.
Table 5: Results of the post-hoc Tukey test for datasets with $Q\geq 6$ (left part) and for datasets with $Q<6$ (right part), in terms of RPS.
@@
As a general comment, by comparing left and right parts of Tables 3, 4 and 5, it is clear that differences are more evident in performance when the number of classes increases. This comes from the fact that the possible costs of misclassification increases, and ordinal splitting criteria can better exploit the ordinality of the target label and its relationship with the input data.
5.2 Statistical analysis for datasets with $Q<6$
Focusing on the MAE metric, the hypothesis of normality is accepted, then, the ANOVA II is applied, showing that there are significant differences between the mean values of the five splitting criteria (p-value$=0.049$). Thus, the post-hoc Tukey test is conducted. The results are shown in the right part of Table 3. It can be observed that OGini achieves the best mean results. However, there are no significant differences between OGini, RI, and WIG with respect to Gini and IG.
In terms of QWK, since the hypothesis of normality is accepted, the ANOVA II test is applied indicating that no significant differences are found with respect to the mean results obtained by the splitting criteria, achieving a p-value$=0.446$. In this case, the post-hoc Tukey test is not performed (in Table 4 all the splitting criteria are in the same group).
Finally, for the RPS metric the hypothesis of normality is accepted again. Then, the ANOVA II test is conducted, showing significant differences in the mean values between the different splitting criteria. Hence, the post-hoc Tukey test is performed. The results (see Table 5) indicate that OGini is the best splitting criteria in mean results, but no significant differences are found with respect to Gini and IG.
5.3 Analysis of the confusion matrices
@@
Table 6: Confusion matrices obtained in the test set by Gini and OGini in the census-2-5 dataset together with the QWK, MAE and RPS values.